# Learning Multimodal Embeddings for Traffic Accident Prediction and Causal Estimation

### v2

@@ 3.3. Causal Estimation Results
Building on the predictive performance of our model, we analyze how contextual factors influence accident risk. We first conduct descriptive analyses by directly aggregating accident counts across different conditions, such as season, road type, and traffic volume, to identify coarse-grained trends. To further validate these observations, we apply causal estimation techniques using learned multimodal embeddings to quantify the impact of specific factors.
@@
Most states show fewer accidents in spring than in winter. This pattern suggests that winter weather conditions contribute to higher accident risk. Snow and ice are likely major factors. In warmer states such as Nevada, where winter conditions are mild, accident counts remain stable across seasons. This pattern is consistent with the causal estimates in Table 6. The ATT values are higher in colder states like Montana and Iowa, and much lower in Nevada. As an example, Figure 3 shows the seasonal accident pattern in Massachusetts, where accident frequency peaks in winter.
@@
The results are shown in Figure 3. Motorways have the highest average accident frequency, and the gap compared to other road types is large. This is reasonable because vehicles on motorways travel at high speeds, which reduces reaction time during unexpected events. Trunk roads show the second highest accident frequency. They also carry fast-moving traffic, which creates similar road accident risks.
@@
In our causal analysis, we treat motorways as the intervention group and obtain an ATT of $21.9\%$. This result indicates that road type has a measurable effect on accident occurrence.
@@
When precipitation is below $40$ mm, accident totals stay relatively low. Once precipitation exceeds $60$ mm, the numbers increase rapidly and reach more than $1.6\times 10^{6}$ at $160$ mm. This pattern reflects the higher risks during heavy rainfall, likely caused by reduced traction, limited visibility, and longer stopping distances. The causal estimation analysis supports this result, showing an average ATT of $24.2\%$.
@@
Our analysis reveals that as traffic volume increases up to approximately $200$ vehicles per day, the number of accidents rises accordingly. However, beyond this threshold, the accident rate begins to decline gradually. This trend is reasonable: higher traffic volumes often lead to congestion, which reduces vehicle speeds and, consequently, lowers the likelihood of severe accidents. To understand the relationship between traffic volume and accident frequency, we visualize the distribution in Figure 5(b).
@@ 4. Related Work
Traffic accident prediction is a dynamic task that has been explored using various foundational models. The critical role of spatial and temporal features in time-series forecasting is well established in prior work (undefay, ). In addition, image-based features have proven valuable in related applications such as road attribute inference (undefaad, ; undefaag, ). We now discuss several areas of research that are most related to our distributions.
@@
GraphGPT (undefaax, ) shows that LLMs can be aligned with graph structure through instruction tuning. UrbanGPT (undefaaw, ) further shows that instruction tuning can align spatio-temporal signals with LLMs, allowing LLMs to handle numeric time-series data and achieve strong zero-shot forecasting across urban tasks. Also, HiGPT (undefaay, ) demonstrates that instruction-tuned LLMs can generalize across heterogeneous graphs by encoding node and edge semantics in natural language. (undefaal, ) analyze the generalization in graph neural networks using PAC-Bayesian bounds. (undefaan, ) use higher-order task affinities to boost multitask learning on graphs. (undefaat, ) use gradient-based estimation to accelerate this task affinity computation. GradEx (undefaau, ) introduces a first-order approach for scalable model fine-tuning. EnsembleLoRA (undefaaae, ) uses GradEx in multitask learning. GradSel (undefaaaj, ) extends to the in-context learning setting and evaluates the generalization ability of language models.
@@
Spatiotemporal mining for traffic prediction.
The importance of jointly modeling spatial and temporal dependencies over graph-structured data for time series prediction is well recognized. Recent advances have focused on developing powerful deep learning architectures to capture these complex patterns from structured data like traffic flow and road graphs. For instance, DCRNN (undefay, ) treats traffic flow as a diffusion process on directed graphs and incorporates this into a sequence-to-sequence recurrent architecture with scheduled sampling. To improve training efficiency, STGCN (undefaz, ) eliminates recurrent units by stacking graph and temporal convolutions into a fully convolutional framework. TEMPO (undefaaq, ) introduces a novel prompt-based generative pre-training framework that integrates time series decomposition with transformer architectures. MG-TAR (undefaap, ) leverages dangerous driving statistics as near-miss indicators within a multi-view GNN for citywide risk prediction.
The advent of high-resolution remote sensing has opened up new possibilities for analyzing the physical state of road networks at scale (undefaaac, ; undefaaz, ; undefaav, ; undefax, ). Satellite imagery provides visual cues that are strongly correlated with accident risk, but are difficult to capture in traditional datasets. Researchers have used this data for a range of safety-related tasks. Some approaches use imagery to assess proxy variables for safety, such as monitoring road surface conditions or automatically detecting infrastructure like pedestrian crossings.
Multimodal road network analysis.
Our approach builds on prior research that combines visual models with graph-based learning for road network analysis. The RoadTagger system (undefaad, ) introduced an end-to-end architecture that combines a CNN with a GNN to infer road attributes, such as lane counts, from satellite imagery. By propagating visual features along the road graph, the GNN overcomes the limited receptive field of a CNN, enabling robust inference in the presence of occlusions. Subsequently, the work on Inferring High-Resolution Traffic Accident Risk Maps (undefaah, ) developed a deep learning framework to fuse multiple data sources, including satellite imagery, GPS trajectories, road maps, and historical accidents, to generate fine-grained (5m resolution) risk maps. This work demonstrated how to overcome data sparsity challenges by leveraging a rich, multimodal context. The challenge of integrating heterogeneous data modalities on a graph is also studied in multimodal graph learning (undefaaj, ). Our work contributes to this area in the context of road safety modeling.
Table 7. A summary of our dataset and several existing datasets. Our dataset combines large-scale traffic accident records with aligned satellite imagery, while prior datasets either lack image data or have limited volume.
@@
METR-LA (undefay, ) $2018$ $6M+$ Traffic Forcast
@@ 5. Discussion
The satellite images in our dataset capture road characteristics that change very slowly, such as lane width, curvature, and intersection density. These features remain steady across long periods, so small differences between the imagery date and the accident records do not create clear sources of bias. The dynamic elements of accident risk come from other modalities. Weather observations and time-varying traffic volume (AADT) provide the signals that describe short-term changes in road conditions.
@@
Seen from this angle, our dataset opens up several directions for future work. It provides a strong benchmark for multimodal learning methods that operate on both graph and image data. It also demonstrates that visual cues, structural features, and traffic signals can work together to reveal fine-grained patterns in accident risk. This creates space for developing stronger fusion models, exploring more detailed temporal patterns, and building tools that help agencies use data-driven insights for safety interventions.
The MMTraCE package. To support future research, we release our dataset on Huggingface along with an easy-to-use Python package for streamlined access. The full dataset is available via the Datasets library and includes road network graphs, accident records, weather data, traffic volume statistics, and satellite image embeddings. The embeddings are generated using Vision Transformer and CLIP models. We also include a set of satellite images with the package for all states.
With a single line of code, users can load the complete dataset. To retrieve data for a specific state, one simply specifies the state name, and the package will automatically download, cache, and return the corresponding dataset object. We also provide functionality to extract accident records and features for any given month. Additionally, the package includes a trainer module for training and evaluating baseline models used in our framework, which encompasses basic GNN models, spatial-temporal models, and fusion models. We add some examples of the usage in Appendix B.3.
@@ 6. Conclusion
We present a large and up-to-date multimodal traffic dataset that includes road networks, weather data, traffic volumes, accident records, and satellite images from six U.S. states. The dataset is designed to support research on traffic accident prediction and to provide a comprehensive view of the factors that shape road safety. It brings together long-term records from multiple public sources and aligns them at the road-segment level, making it suitable for both predictive modeling and causal analysis.
@@ Acknowledgments
Thanks to Yan Liu for insightful discussions related to this project. We are also grateful to Abhinav Nippani and Dongyue Li for their helpful feedback during the early stages of this work. We appreciate the anonymous referees for their constructive comments and suggestions. This project is supported by Northeastern University’s Transforming Interdisciplinary Experiential Research (TIER) $1$ program. The work of Z. Zhang, M. Duan, and H. R. Zhang is additionally supported in part by NSF award IIS-2412008. The views and opinions expressed in this paper are solely those of the authors and do not necessarily reflect those of the NSF.
@@ Appendix A Data Collection Procedure
In this section, we provide more details about our data collection procedure, including how to construct the road network, how to align the location of accident records, how to normalize traffic volumes from different formats, and how to obtain weather features. Then, we summarize all the features we collected.
@@ A.1. Road Networks
From the raw data, we extract a list of nodes and edges. For each node, we retain the node ID along with its geographic coordinates (latitude and longitude). For each edge, we record the start node ID, end node ID, one-way or not, the corresponding road type, and the length of the road.
@@ A.2. Traffic Accident Records
We collect the traffic accident data provided by each state’s DOT, which varies in different formats.
In our analysis, we focus specifically on extracting the latitude and longitude information from these records. When geographic coordinates are not directly available in the original dataset, we use the textual address descriptions, such as street names or intersections, and query the corresponding latitude and longitude using the Google Maps API. This step ensures that all accident locations are consistently represented in a geospatial format, which is critical for downstream spatial analysis and visualization.
@@ A.3. Traffic Volume
Some states provide clean CSV exports containing longitude and latitude fields, while others (such as Massachusetts and Montana) publish data in more complex formats, including KML and GeoJSON, with inconsistent column names and nested spatial structures. In these cases, we developed custom parsers for each state and manually mapped the relevant fields to extract usable geospatial information. Additionally, the temporal resolution differs by state, and in some cases, the data spans dozens of measurement points per segment, which requires iterative merging and normalization. Often, the same road segment is represented by multiple sensor points along its path, with no explicit intersection or highway labeling, so we heuristically aggregated these measurements to associate one traffic volume value with each edge.
We report a distribution map of accidents and AADT records in Figure 8(b). It indicates that the traffic volume monitor has covered nearly all potential road segments with accidents.
@@ A.4. Weather Statistics
The weather data is extracted using the Meteostat API, focusing on monthly weather statistics. For each node in the road network, the nearest weather station is identified based on geographic coordinates (latitude and longitude). We then retrieve monthly weather data for each location over a specified time range. The time range aligns with the time range of the traffic accident record. After fetching the data, it merges this with the corresponding node ID and coordinates, and stores the results in CSV files. Finally, all files are concatenated to create a comprehensive dataset matching historical weather conditions to each road network node.
@@
In summary, we report all the features we collect from our road networks in this section. For node features, we collect $9$ types of features: latitude, longitude, and satellite image for static features, average surface temperature, max surface temperature, min surface temperature, total precipitation, average wind speed, and sea level air pressure for dynamic features. For edge features, we collect $3$ types of features: location, length, and road type for static features, and traffic volume (annual average daily traffic) for dynamic features.
@@ B.1. Implementation Details
We report the hyperparameters in our experiments and give the formal definition of the fusion methods and evaluation metrics in the main result.
@@
Finally, we implement the mixture of experts (MoE) approach, which leverages multiple specialized networks, referred to as experts, that each learn to combine features from different perspectives. Given the visual feature $z$ and the graph embedding $x^{(\text{GNN})}$, we first construct a shared representation by concatenating the two modalities: $\tilde{x}=[x^{(\text{GNN})}\parallel z^{\prime}]$. This shared feature vector is then fed into a set of $K$ expert networks, each parameterized by its own learnable weights. Each expert $f_{k}$ produces an output representation:
@@
To dynamically select and combine the outputs of these experts, we utilize the gated network that computes a probability distribution over experts for each node. This gating mechanism is realized as a small multilayer perceptron followed by a softmax activation:
@@
This MoE architecture enables the model to dynamically select among multiple fusion pathways, capturing complex relationships between different modalities. With a learned gating mechanism, the MoE framework can adaptively choose the most relevant experts for each node, resulting in richer, more flexible multimodal representations.
@@
For all baselines, we construct edge representations by concatenating the embeddings of the two connected nodes with the associated edge features. Node embeddings are fixed to a dimension of $128$. We use two-layer MLP and GNN architectures, both with a hidden size of $256$. The GNN, MLP, and fusion modules are fully trainable. All models are optimized using the Adam algorithm with a learning rate of $0.001$ and trained for $30$ epochs.
We split all the accident records into training, validation, and test sets for each state based on different periods of years. The summarization is shown in Table 8. Since our current analysis is at the monthly level, we note that our framework can also be utilized for conducting analysis at the yearly level.
@@
Beyond direct matching, we also estimate causal effects with propensity score-based adjustments (PSM). We model the treatment assignment probability with

### v3

@@ 3.3. Causal Estimation Results
Building on our model’s predictive performance, we analyze how contextual factors influence accident risk. We first conduct descriptive analyses by directly aggregating accident counts across conditions such as season, road type, and traffic volume to identify coarse-grained trends. To further validate these observations, we apply causal estimation techniques using learned multimodal embeddings to quantify the impact of specific factors.
@@
Most states show fewer accidents in spring than in winter. This pattern suggests that winter weather conditions increase the risk of accidents. Snow and ice are likely major factors. In warmer states such as Nevada, where winter conditions are mild, accident counts remain stable across seasons. This pattern is consistent with the causal estimates in Table 6. We set winter as the treatment group and spring as the control group. The ATT values are higher in colder states such as Montana and Iowa, and much lower in Nevada. As an example, Figure 3 shows the seasonal accident pattern in Massachusetts, where accident frequency peaks in winter.
@@
The results are shown in Figure 3. Motorways have the highest average accident frequency, and the gap compared to other road types is large. This is reasonable because vehicles on motorways travel at high speeds, reducing reaction time in case of unexpected events. Trunk roads show the second-highest accident frequency. They also carry fast-moving traffic, which poses similar risks of road accidents.
@@
In our causal analysis, we treat motorways as the treatment group and other road types as the control group. The estimated ATT is $21.9\%$ shown in Table 6. This suggests that motorway road segments contribute to a higher likelihood of accident occurrence after controlling for other factors.
@@
When precipitation is below $40$ mm, accident totals remain relatively low. As precipitation increases beyond $60$ mm, accident counts rise rapidly and reach more than $1.6\times 10^{6}$ at $160$ mm. This trend suggests that heavy rainfall increases accident risk, likely due to reduced tire traction, lower visibility, and longer stopping distances. To further quantify this effect, we treat precipitation above $120$ mm as the treatment group and precipitation below $40$ mm as the control group. The causal estimation result in Table 6 shows an average ATT of $24.2\%$, supporting the positive effect of high precipitation on accident occurrence.
@@
Our analysis reveals that as traffic volume increases up to approximately $200$ vehicles per day, the number of accidents rises accordingly. However, beyond this threshold, the accident rate declines gradually. This trend is reasonable: higher traffic volumes often lead to congestion, which reduces vehicle speeds and, consequently, lowers the likelihood of severe accidents. To understand the relationship between traffic volume and accident frequency, we visualize the distribution in Figure 5(b).
@@ 4. Related Work
Traffic accident prediction is a dynamic task that has been explored using various foundational models. The critical role of spatial and temporal features in time-series forecasting is well established in prior work (undefay, ). In addition, image-based features have proven valuable in related applications such as road attribute inference (undefaad, ; undefaag, ). We now discuss several areas of research most closely related to our contributions.
@@
GraphGPT (undefaax, ) shows that LLMs can be aligned with graph structure through instruction tuning. UrbanGPT (undefaaw, ) further shows that instruction tuning can align spatio-temporal signals with LLMs, allowing LLMs to handle numeric time-series data and achieve strong zero-shot forecasting across urban tasks. Also, HiGPT (undefaay, ) demonstrates that instruction-tuned LLMs can generalize across heterogeneous graphs by encoding node and edge semantics in natural language. (undefaal, ) analyze the generalization in graph neural networks using PAC-Bayesian bounds. (undefaan, ) use higher-order task affinities to boost multitask learning on graphs. (undefaat, ) use gradient-based estimation to accelerate the computation of task affinity. GradEx (undefaau, ) introduces a first-order approach for scalable model fine-tuning. EnsembleLoRA (undefaaae, ) uses GradEx in multitask learning. GradSel (undefaaaj, ) extends to the in-context learning setting and evaluates the generalization ability of language models.
@@
Table 7. A summary of our dataset and several existing datasets. Our dataset combines large-scale traffic accident records with aligned satellite imagery, whereas prior datasets either lack imagery or have limited volume.
@@
METR-LA (undefay, ) $2018$ $6M+$ Traffic Forecast
@@
Spatiotemporal mining for traffic prediction.
The importance of jointly modeling spatial and temporal dependencies over graph-structured data for time series prediction is well recognized. Recent advances have focused on developing powerful deep learning architectures to capture these complex patterns in structured data such as traffic flow and road graphs. For instance, DCRNN (undefay, ) treats traffic flow as a diffusion process on directed graphs and incorporates this into a sequence-to-sequence recurrent architecture with scheduled sampling. To improve training efficiency, STGCN (undefaz, ) eliminates recurrent units by stacking graph and temporal convolutions into a fully convolutional framework. TEMPO (undefaaq, ) introduces a novel prompt-based generative pre-training framework that integrates time series decomposition with transformer architectures. MG-TAR (undefaap, ) leverages dangerous driving statistics as near-miss indicators within a multi-view GNN for citywide risk prediction.
The advent of high-resolution remote sensing has opened up new possibilities for analyzing the physical state of road networks at scale (undefaaac, ; undefaaz, ; undefaav, ; undefax, ). Satellite imagery provides visual cues that are strongly correlated with accident risk, but are difficult to capture in traditional datasets. Researchers have used this data for a range of safety-related tasks. Some approaches use imagery to assess safety proxies, such as monitoring road surface conditions or automatically detecting infrastructure like pedestrian crossings.
Multimodal road network analysis.
Our approach builds on prior research that combines visual models with graph-based learning for road network analysis. The RoadTagger system (undefaad, ) introduced an end-to-end architecture that combines a CNN with a GNN to infer road attributes, such as lane counts, from satellite imagery. By propagating visual features along the road graph, the GNN overcomes the limited receptive field of a CNN, enabling robust inference in the presence of occlusions. Subsequently, the work on Inferring High-Resolution Traffic Accident Risk Maps (undefaah, ) developed a deep learning framework that fuses multiple data sources, including satellite imagery, GPS trajectories, road maps, and historical accident data, to generate fine-grained (5m resolution) risk maps. This work demonstrated how to overcome data sparsity challenges by leveraging a rich, multimodal context. The challenge of integrating heterogeneous data modalities on a graph is also studied in multimodal graph learning (undefaaj, ). Our work contributes to this area in the context of road safety modeling.
@@ 5. Discussion
The satellite images in our dataset capture road characteristics that change very slowly, such as lane width, curvature, and intersection density. These features remain stable over long periods, so small differences between the imagery date and the accident records do not introduce clear sources of bias. The dynamic elements of accident risk come from other modalities. Weather observations and time-varying traffic volume (AADT) provide signals describing short-term changes in road conditions.
@@
Seen from this angle, our dataset opens up several directions for future work. It provides a strong benchmark for multimodal learning methods that operate on both graph and image data. It also demonstrates that visual cues, structural features, and traffic signals can work together to reveal fine-grained patterns in accident risk. This creates space to develop stronger fusion models, explore more detailed temporal patterns, and build tools that help agencies use data-driven insights for safety interventions.
Our causal analysis relies on the common assumption that treatment and control units are independent. In road networks, this assumption may not hold exactly, since neighboring nodes and edges are spatially connected and can influence each other. Therefore, the estimated effects should be interpreted as approximate associations under this modeling assumption.
The MMTraCE package. To support future research, we release our dataset on Huggingface along with an easy-to-use Python package for streamlined access. The full dataset is available via the Datasets library and includes road network graphs, accident records, weather data, traffic volume statistics, and satellite image embeddings. The embeddings are generated using Vision Transformer and CLIP models. We also include a set of satellite images for all states in the package.
With a single line of code, users can load the complete dataset. To retrieve data for a specific state, one simply specifies the state name, and the package will automatically download, cache, and return the corresponding dataset object. We also provide functionality to extract accident records and features for any given month. Additionally, the package includes a trainer module for training and evaluating baseline models in our framework, encompassing basic GNN models, spatiotemporal models, and fusion models. We add some examples of the usage in Appendix B.3.
@@ 6. Conclusion
We present a large, up-to-date multimodal traffic dataset comprising road networks, weather data, traffic volumes, accident records, and satellite imagery from six U.S. states. The dataset is designed to support research on traffic accident prediction and to provide a comprehensive view of the factors that shape road safety. It brings together long-term records from multiple public sources and aligns them at the road-segment level, making it suitable for both predictive modeling and causal analysis.
@@ Acknowledgments
Thanks to Yan Liu, Abhinav Nippani, Dongyue Li, and Ruoxuan Xiong for discussions and feedback on this project. We appreciate the anonymous referees for their comments and suggestions.
This project is supported by Northeastern University’s Transforming Interdisciplinary Experiential Research (TIER) $1$ program. The work of Z. Zhang, M. Duan, and H. R. Zhang is additionally supported in part by NSF award IIS-2412008. The views and opinions expressed in this paper are solely those of the authors and do not necessarily reflect those of the National Science Foundation.
@@ Appendix A Data Collection Procedure
In this section, we provide more details on our data collection procedure, including how to construct the road network, align the locations of accident records, normalize traffic volumes across different formats, and obtain weather features. Then, we summarize all the features we collected.
@@ A.1. Road Networks
From the raw data, we extract a list of nodes and edges. For each node, we retain the node ID along with its geographic coordinates (latitude and longitude). For each edge, we record the start node ID, end node ID, whether it is one-way, the corresponding road type, and the road length.
@@ A.2. Traffic Accident Records
We collect traffic accident data from each state’s DOT, which is provided in different formats.
In our analysis, we focus on extracting latitude and longitude from these records. When geographic coordinates are not directly available in the original dataset, we use the textual address descriptions, such as street names or intersections, and query the corresponding latitude and longitude using the Google Maps API. This step ensures that all accident locations are consistently represented in a geospatial format, which is critical for downstream spatial analysis and visualization.
@@ A.3. Traffic Volume
Some states provide clean CSV exports containing longitude and latitude fields, while others (such as Massachusetts and Montana) publish data in more complex formats, including KML and GeoJSON, with inconsistent column names and nested spatial structures. In these cases, we developed custom parsers for each state and manually mapped the relevant fields to extract usable geospatial information. Additionally, the temporal resolution varies by state, and in some cases, the data spans dozens of measurement points per segment, requiring iterative merging and normalization. Often, the same road segment is represented by multiple sensor points along its path, with no explicit intersection or highway labeling, so we heuristically aggregated these measurements to assign a single traffic volume value to each edge.
We present a distribution map of accident counts and AADT records in Figure 8(b). It indicates that the traffic volume monitor has covered nearly all potential road segments that have had accidents.
@@ A.4. Weather Statistics
The weather data is extracted using the Meteostat API, focusing on monthly weather statistics. For each node in the road network, the nearest weather station is identified based on geographic coordinates (latitude and longitude). We then retrieve monthly weather data for each location over a specified time range. The time range aligns with that of the traffic accident record. After fetching the data, it merges it with the corresponding node IDs and coordinates and stores the results in CSV files. Finally, all files are concatenated to create a comprehensive dataset matching historical weather conditions to each road network node.
@@
In summary, we report all features collected from our road networks in this section. For node features, we collect $9$ types: static features (latitude, longitude, and satellite imagery) and dynamic features (average surface temperature, maximum surface temperature, minimum surface temperature, total precipitation, average wind speed, and sea-level air pressure). For edge features, we collect $3$ types of static features (location, length, and road type) and $1$ dynamic feature (annual average daily traffic).
@@ B.1. Implementation Details
We report the hyperparameters used in our experiments and provide the formal definitions of the fusion methods and evaluation metrics in the main result.
@@
Finally, we implement the mixture-of-experts (MoE) architecture, which leverages multiple specialized networks, referred to as experts, that each learn to combine features from different perspectives. Given the visual feature $z$ and the graph embedding $x^{(\text{GNN})}$, we first construct a shared representation by concatenating the two modalities: $\tilde{x}=[x^{(\text{GNN})}\parallel z^{\prime}]$. This shared feature vector is then fed into a set of $K$ expert networks, each parameterized by its own learnable weights. Each expert $f_{k}$ produces an output representation:
@@
To dynamically select and combine the outputs of these experts, we use a gated network that computes a probability distribution over the experts for each node. This gating mechanism is realized as a small multilayer perceptron followed by a softmax activation:
@@
This MoE architecture enables the model to dynamically select among multiple fusion pathways, capturing complex relationships between different modalities. With a learned gating mechanism, the MoE framework can adaptively select the most relevant experts for each node, yielding richer, more flexible multimodal representations.
@@
For all baselines, we construct edge representations by concatenating the embeddings of the two connected nodes with the associated edge features. Node embeddings are fixed to a dimension of $128$. We use two-layer MLP and GNN architectures, both with a hidden size of $256$. The GNN, MLP, and fusion modules are fully trainable. All models are optimized with the Adam optimizer at a learning rate of $0.001$ and trained for $30$ epochs.
We split all the accident records into training, validation, and test sets for each state, based on different time periods. The summarization is shown in Table 8. Since our current analysis is at the monthly level, we note that our framework can also be used to conduct analysis at the yearly level.
@@
Beyond direct matching, we also estimate causal effects using propensity score matching (PSM). We model the treatment assignment probability with